OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (1, -9) and (3, -1).

hero (hero):

Insert each point one at a time into the formula y = mx + b creating two equations: -9 = m + b -1 = 3m + b Next isolate b in each equation: b = -9 - m b = -1 - 3m Divide both sides of the first equation by -1 to get -b = 9 + m b = -1 - 3m Combine both equations together to get 0 = 8 -2m Add 2m to both sides to get 2m = 8 Finish solving for m. Let me know what you get.

hero (hero):

You have to solve for m and b to get the slope and y-intercept.

OpenStudy (anonymous):

2m+8=10 or do i have to *

hero (hero):

:/

hero (hero):

Do you know how to isolate a variable in an equation?

OpenStudy (anonymous):

no

hero (hero):

I'm sorry to hear that. That's basic algebra. You should know how to do it.

OpenStudy (anonymous):

you have to find slope first, which you can do when given 2 points. Do you know that formula?

hero (hero):

You're supposed to divide both sides by 2 to get m = 8/2 m = 4

hero (hero):

m = the slope of the line you're trying to find

hero (hero):

Now if you substitute m back into one of the original equations above, you can get b: b = -9 - m b = -9 - 4 @amber517 can you solve for b here?

OpenStudy (anonymous):

ok from what i was look at top i massed up and thought you + 2m to 8 and

hero (hero):

Can you solve for b?

hero (hero):

Tell me what b equals

hero (hero):

:/

OpenStudy (anonymous):

would the b=1 and b=3

hero (hero):

Did you even read anything I posted? :/

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Amber - you can find slope by finding the change in y divided by the change in x. Once you find the slope (m) you can then put it in the slope-intercept form. y-y1=m(x-x1). Which is what they are asking for. =)

hero (hero):

@mmencer, we were getting there. All @amber517 had to do was find m and b The slope intercept form is obviously y = mx + b

hero (hero):

@mmencer, what you posted is called the point-slope form

hero (hero):

@mmencer, you're confused.

OpenStudy (anonymous):

i'm confused

OpenStudy (anonymous):

I'll leave you two to finish...sorry to interrupt. Good luck, Amber.

OpenStudy (anonymous):

so i'm tryin to understand this i'm dum when it comes to algebra

OpenStudy (anonymous):

-9 = m + b -1 = 3m + b when you did this what happen to the other 1

OpenStudy (anonymous):

shouldn't it be -9=1m+b -1=3m+b

hero (hero):

m = 1m It's the same thing

hero (hero):

Remember, anything times 1 is itself

hero (hero):

1 x 9 = 9 1 x m = m

OpenStudy (anonymous):

ok ok would it be -13

OpenStudy (anonymous):

b=-13

OpenStudy (anonymous):

m=(y2-y1)/(x2-x1) m=(-1--9)/(3-1) m=8/2 m=4 y=mx+b -1=4(3)+b -1=12+b -13=b y=4x-13

hero (hero):

Yes, finally

hero (hero):

Well, you didn't do it the way I wanted you to do it, but you did it.